# MASS: Motion-Aware Spatial–temporal Grounding for Physics Reasoning and Comprehension in Vision-Language Models

### v1

@@ Abstract
Vision Language Models (VLMs) perform well on standard video tasks but struggle with physics-driven reasoning involving motion dynamics and spatial interactions. This limitation reduces their ability to interpret real or AI-Generated Content (AIGC) videos and to generate physically consistent content. We present a novel approach to address this gap by translating physical-world context cues into interpretable representations aligned with VLMs’ perception, comprehension, and reasoning. We introduce a comprehensive benchmark, MASS-Bench, consisting of $4,350$ real-world and AIGC videos and $8,361$ free-form video question–answering pairs focused on physics-related comprehension tasks, with detailed annotations including visual detections and grounding over sub-segments, as well as full-sequence 3D motion tracking of entities. We further present MASS, a model-agnostic approach that injects spatial–temporal signals into the VLM language space via depth-based 3D encoding and visual grounding, coupled with a motion tracker for object dynamics. To strengthen cross-modal alignment and reasoning, we apply reinforcement fine-tuning. Experiments and ablations show that our refined VLMs outperform comparable and larger baselines, and prior state-of-the-art models, by $8.7\%$ and $6.0\%$, achieving performance comparable to close-source SoTA VLMs like Gemini-2.5-Flash on physics reasoning and comprehension, validating the effectiveness of our approach.
@@ 1 Introduction
Vision Language Models (VLMs) demonstrate strong reasoning and comprehension in standard video tasks such as captioning [li2024wolf], event recognition [li2025eventvl] and scene understanding [yu2025evaluating]. However, they struggle with complex visual cues that involve intertwined 3D spatial layouts [cai2025holisticevaluationmultimodalllms, yang2025thinking], motion patterns [chen2024motionllm], and temporal dynamics [zhou2025vlm4d, fu2025video]. Achieving robust physical understanding requires VLMs not only to perceive visual cues, but also to internalize real-world physical principles and commonsense expectations about object behavior [chow2025physbench]. Such physical reasoning, central to human-level video understanding, remains challenging, as models must connect visual evidence with underlying physical dynamics and reason about whether observed events align with or violate real-world physics.
A central difficulty lies in the implicit nature of physical laws. Unlike tasks such as segmentation or object grounding, where the supervision is explicit and localized, real-world physics must be inferred from indirect and often ambiguous visual evidence. Many phenomena governed by the same underlying principle can manifest in drastically different visual forms. For example, an apple falling and a person standing still are both influenced by gravity but differ substantially in spatial configuration, motion patterns, and temporal structure. As a result, VLMs often fail to generalize across diverse physical processes, a challenge further magnified by the scarcity of datasets with dense spatiotemporal and motion-level annotations. Without such supervision, models tend to memorize superficial correlations rather than develop physics-grounded reasoning. Effective physics comprehension requires several demanding prerequisites, spatial and temporal understanding, motion tracking, object detection, and visual grounding, yet these components are rarely annotated in sufficient detail. This lack of rich spatiotemporal data widens the gap between physics-driven video dynamics and VLM cognition, ultimately degrading their ability to reason coherently about real-world physical behavior.
@@
Compounding these issues, existing VLMs are typically trained on large collections of real-world videos, where object motions naturally conform to physical laws. While such data can help models implicitly acquire certain motion patterns, it also encourages strong language and visual priors: VLMs tend to assume that observed actions are physically plausible simply because they resemble frequent patterns in the training set. This reliance on priors becomes particularly problematic in the era of AI-generated (AIGC) videos, where implausible trajectories, inconsistent depth cues, or temporally incoherent motions are common. Prior work has shown that VLMs frequently hallucinate or overlook physics abnormalities [bansal2025videophy, li2025videohallu]; for instance, a model may conclude that an apple must be falling simply because it appears near a tree, reflecting prior expectations rather than actual visual evidence.
These challenges highlight the need for methods that go beyond raw video pixels and high-level textual supervision. Instead of training VLMs from scratch, which is infeasible and very costly in time and computational resource, one more data-efficient way is to augment VLMs with structured spatial and temporal representations that capture fine-grained object motions, interactions, and geometry. Recent advances in motion grounding [deng2025motion], visual grounding, detailed temporal modeling [ye2025re], and action analysis [du2025motionsight, hong2025motionbench] demonstrate that smaller specialized models can provide highly accurate spatial–temporal cues. However, existing approaches rarely explore how such modules can be systematically integrated into VLMs to elevate higher-level physics reasoning. Meanwhile, prior physics-oriented datasets [li2025videohallu, li2025worldmodelbench, bansal2025videophy] primarily rely on coarse annotations, limiting their capacity to induce deep, mechanism-level understanding.
@@ 2 Related Work
Physical Reasoning and Abnormality Detection in Videos: With the rise of AIGC video generation, abnormalities such as prompt misalignment and violations of physics or commonsense remain prevalent across models from LaVIE [wang2025lavie], SORA [brooks2024video], and CogVideoX [yang2024cogvideox] to newer systems like VEO3 [veo3], Wan2.2 [wan2025], and COSMOS [agarwal2025cosmos]. Detecting and reducing these issues has become a major area of research. Early work such as VideoScore [he2024videoscore] trained synthetic evaluators aligned with human judgment to assess abnormalities, while large benchmarks [li2025worldmodelbench, qin2024worldsimbench, motamed2025generative, chow2025physbench, li2025videohallu, zhang2025morpheus] provide annotated data for physics and commonsense reasoning. WorldScore [duan2025worldscore] offers unified evaluation across controllability, quality, and dynamics, and CRAVE [sun2025content] supports content-rich assessment using text–temporal fusion and motion-fidelity modeling. Recent efforts aim at mitigation via prompt engineering or external physical cues: [kang2024far] uses physics-based features, PhyT2V [xue2025phyt2v] applies chain-of-thought reasoning, and VideoPhy-2 [bansal2025videophy] provides action-centric evaluation with corresponding fixes. TRAVL [motamed2025travl] enhances motion-aware physical plausibility judgment, offering a unified framework for improving physical realism. Despite this progress, physical and commonsense abnormalities remain challenging, and our approach addresses this persistent gap.
Spatial, Temporal and Motion Understadning in VLM: As VLMs evolve, their limitations in video-based perception have inspired more research in terms of spatial and temporal reasoning. For spatial understanding, SpatialRGPT [cheng2024spatialrgpt] enhances 3D perception through regional representations and depth cues; VG-LLM [zheng2025learning] extracts 3D geometry directly from videos; and LayoutVLM [sun2025layoutvlm] uses semantic priors with differentiable layout optimization to generate plausible 3D arrangements. For temporal reasoning, SlowFast-LLaVA [xu2024slowfast] captures detailed spatial and long-range temporal cues through a slow–fast design, while TVS [fan2025video] improves alignment between queries and focus-critical segments. Long-video search is further advanced by T* [ye2025re], which reframes temporal retrieval as spatial search on LV-Haystack. Other works address action and motion understanding: OpenMixer [bao2025exploiting] enables open-vocabulary action detection; Video-MME [hong2025motionbench] exposes weaknesses in long-video comprehension; MotionSight [du2025motionsight] introduces zero-shot prompts for fine-grained motion perception; and GroundMoRe [deng2025motion] supports motion-grounded reasoning via spatiotemporal masks. Despite these advances, most methods do not explicitly address physics-centric reasoning, which requires jointly integrating spatial, temporal, and motion understanding with physics-based commonsense.
Video Understanding in VLM: State-of-the-art VLMs like Qwen-VL [bai2025qwen2] and InternVL [zhu2025internvl3] exhibit strong general video understanding capabilities, but struggle in terms of complex spatial, temporal, and long-video reasoning [li2025survey]. LongVLM [weng2024longvlm] enhances long video comprehension via hierarchical token merging, while VideoMind [liu2025videomind] uses a role-based agent with Chain-of-LoRA for efficient temporal reasoning. Video-Holmes [cheng2025video] benchmarks complex reasoning from suspense films, showing VLMs can still miss multi-segment clues. EgoLife [yang2025egolife] supports egocentric video assistants with long-context retrieval. To mitigate hallucinations, MASH-VLM [bae2025mash] disentangles spatial-temporal attention, and MVoT [li2025imagine] enables visual reasoning traces for spatial reasoning. Scaling the data also helps: [chen2024spatialvlm] builds a 2B-sample dataset for 3D spatial VQA and robotics. Training strategies further advance reasoning. Video-R1 [feng2025video] applies R1-style reinforcement learning with T-GRPO, surpassing GPT-4o on VSI-Bench, while VideoChat-R1 [li2025videochat] uses reinforcement fine-tuning to boost spatio-temporal tasks like grounding and tracking.
@@ 3 MASS-Bench: A Motion-Grounded Physics Reasoning and Comprehension Benchmark
Figure 2: Data Exhibition of MASS-Bench. MASS-Bench provides two question types—factual and critical-thinking—to evaluate physics-driven video understanding. For each video–question–answer pair, we supply rich motion-grounding annotations, including temporal segmentation, entity-level visual grounding, temporal profiles across the full video, and motion attributes such as first/last positions and 3D displacement vectors. These structured spatial–temporal cues transform complex physics-related perception into interpretable representations that support more reliable physical reasoning. Additional dataset details are provided in the Appendix A.
Data Collection: MASS-Bench is a free-form video question–answering dataset aimed at strengthening VLMs’ physics understanding, built from real-world videos in Motion-Sight [du2025motionsight] and ActivityNet [caba2015activitynet], as well as AIGC videos from VideoPhy2 [bansal2025videophy], VideoHallu [li2025videohallu], and models such as COSMOS [agarwal2025cosmos], Wan2.2 [wan2025], and Sora [brooks2024video]. MASS-Bench contains $4,350$ videos and $8,361$ free-form QA pairs covering diverse physics-related reasoning tasks. Each example includes positive and negative demonstrations with detailed human annotations, enabling models to learn correct physical processes from real videos while identifying artifacts and violations commonly found in AIGC content.
@@
Categorization: To capture the spectrum of physics-related cognition required in MASS-Bench, we group all questions into five categories reflecting the underlying visual and physical reasoning processes:
Category Count Pct.
SU (Spatial Understanding) 2,785 33.31%
TU (Temporal Understanding) 1,633 19.53%
PA (Physical Abnormality Detection) 1,432 17.13%
PC (Physics Comprehension) 1,304 15.60%
MAR (Motion & Action Recognition) 1,205 14.41%
Table 1: Distribution of question categories in MASS-Bench. We group all questions into five categories that reflect the underlying visual and physics-driven reasoning processes, based on the type of physical phenomena probed and the level of reasoning required.
We classify these categories from easy to challenging based on the depth of reasoning involved. VLMs must first establish spatial and temporal understanding, then recognize motion patterns, before progressing to higher-level physics comprehension and detecting real-world violations or abnormalities. The breakdown of the questions among different categories is highlighted in Table 1.
Motion Grounding. For each video–question–answer pair, we provide comprehensive motion-grounding annotations in addition to the question, video, and ground-truth answer. These include: (1) temporal video segmentation indexed by frame ranges; (2) visual grounding for each queried entity specified by the entity name and its bounding box; and (3) an entity-level temporal profile that tracks each grounded entity across the entire video—persisting once detected and left blank in segments where the entity is absent. Within each video segment, we further provide (4) motion-grounding attributes, including the entity’s first and last observed positions to capture coarse spatial layout, and (5) 3D motion vectors representing the entity’s temporal displacement. By encoding these spatial–temporal cues, the dataset converts physics-intensive perceptual challenges into structured textual and mathematical representations, enabling more reliable physical reasoning in multimodal models.
Figure 3: Overview of MASS: We use a model-agnostic approach to enhance visual recognition with explicit spatial and motion awareness. Beyond standard visual transformer encoders that process video inputs (e.g., LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2]), we introduce a visual grounding module to strengthen correlations between queried entities and corresponding visual cues. Depth estimation captures spatial geometry, while motion tracking encodes temporal dynamics across frames. These spatial and temporal signals are fused into motion traces for each entity and tokenized with grounding and temporal features to align them with the language domain. During post-training, we freeze the spatial–temporal encoders and apply reinforcement fine-tuning (RFT) to improve the LLM backbone’s comprehension of the additional multimodal information.
Data Metadata. Our dataset has $6093$ examples for training and $2268$ examples for testing. Our dataset covers videos with average $545.8$ frames per video, corresponding to approximately $19.62$ seconds at an average framerate of $27.37$ FPS. Frame resolution averages $1120\times 702$ pixels across all videos.
@@ 4 MASS: Model-Agnostic Approach
High-quality human annotations are essential for strengthening VLMs’ understanding of physics-related visual contexts, yet data alone cannot address their core limitations in modeling spatial layouts and motion dynamics. Effective physical reasoning requires isolating relevant cues from noisy video content. While one solution is to retrain models with heavy preprocessing pipelines, this incurs substantial cost. Instead, we propose a more data-efficient alternative: augmenting VLMs with dedicated spatial and motion encoders that explicitly extract and represent key visual signals. As shown in Figure 3, our model-agnostic design—motivated by dataset insights and observed VLM weaknesses, combines lightweight architectural refinement with targeted training to enhance spatiotemporal perception, cross-modal alignment, and physics-aware reasoning, ultimately reducing hallucinations and improving physical comprehension.
@@ 4.1 Entity-Centric Visual Grounding
Understanding physics-related dynamics in videos typically requires identifying and tracking one or more specific entities throughout the sequence. However, videos contain rich, high-dimensional spatial–temporal context, making it challenging to isolate and follow these entities across frames. Besides, language queries are often ambiguous about which entities they refer to. Grounding is therefore essential for isolating the correct objects and preserving their spatial–temporal dynamics, which would otherwise be lost or distorted by high-level VLM encoders. Establishing video-scale grounding profiles for the queried entities helps align modalities, filter out irrelevant or distracting information, and mitigate the information loss introduced by LLaVA-style VLMs [liu2023visual], which compress visual signals into limited textual embeddings.
We begin by semantically extracting the key components from each probing question—those most relevant to physics, motion, or other spatial–temporal dynamics—to narrow the target for visual grounding. Using Grounding-DINO [liu2024grounding], we detect bounding boxes for the queried entities and apply a dynamic temporal-resolution scheme that automatically selects an appropriate sub-sequence length based on the input video’s duration. This segmentation strategy increases temporal coherence within each chunk, improving object tracking and downstream processing while balancing accuracy and efficiency. It also preserves grounding consistency over time and reveals temporal artifacts such as sudden appearance or disappearance, which are common in AIGC videos. For each detected entity, we construct a time-aligned grounding profile. Finally, we apply SAM2 [ravi2024sam, ren2024grounded] to generate segmentation masks for the grounded regions, enabling subsequent spatial–temporal analysis.

### v2

@@ Abstract
Vision Language Models (VLMs) perform well on standard video tasks but struggle with physics-related reasoning that involves motion dynamics and spatial interactions. We present a novel approach to address this gap by translating physical-world context cues into interpretable representations aligned with VLMs’ perception, comprehension, and reasoning. We present MASS, a model-agnostic approach that injects spatial–temporal signals into the VLM language space via depth-based 3D encoding and visual grounding, coupled with a motion tracker for object dynamics. We also contribute a comprehensive benchmark, MASS-Bench, consisting of $4,350$ real-world and AIGC videos and $8,361$ free-form video question–answering pairs focused on physics-related comprehension tasks, with detailed annotations including visual detections and grounding over sub-segments, as well as full-sequence 3D motion tracking of entities. To strengthen cross-modal alignment and reasoning, we apply reinforcement fine-tuning to MASS. Experiments and ablations show that our refined VLMs outperform comparable and baselines of larger size, and prior state-of-the-art (SoTA) models, achieving performance comparable to close-source SoTA VLMs ($2\%$ difference) like Gemini-2.5-Flash on physics reasoning and comprehension.
@@ 1 Introduction
Vision Language Models (VLMs) demonstrate strong reasoning and comprehension in standard video tasks such as captioning [27], event recognition [31] and scene understanding [65]. However, they struggle with complex visual cues that involve intertwined 3D spatial layouts [9, 60], motion patterns [11], and temporal dynamics [69, 21]. Achieving robust physical understanding requires VLMs not only to perceive visual cues, but also to internalize real-world physical principles and commonsense expectations about object behavior [15]. Such physical reasoning, central to human-level video understanding, remains challenging, as models must connect visual evidence with underlying physical dynamics and reason about whether observed events align with or violate real-world physics.
Unlike tasks such as segmentation or object grounding, where the supervision is explicit and localized, real-world physics must be inferred from indirect and often ambiguous visual evidence. For example, an apple falling and a person standing still are both influenced by gravity but differ substantially in spatial configuration, motion patterns, and temporal structure. As a result, VLMs often fail to generalize across diverse physical processes [4, 36], a challenge further magnified by the scarcity of datasets with dense spatiotemporal and motion-level annotations [16, 38]. Without explicit supervision, models tend to memorize superficial correlations rather than develop physics-grounded reasoning [14]. Effective physics comprehension requires several demanding prerequisites, spatial and temporal understanding, motion tracking, object detection, and visual grounding, yet these components are rarely annotated in sufficient detail [33, 53].
@@
Most VLMs are trained primarily on real-world videos, where object motion usually follows physical laws. Although such data helps models capture common motion patterns, it also reinforces strong visual and language priors: models often assume that observed events are physically plausible because they match frequent patterns in the training data. This becomes more problematic with AIGC videos, where implausible trajectories, inconsistent depth, and temporally incoherent motion are common. Prior studies show that VLMs often miss or hallucinate such physics violations [5, 36]. For example, a model may infer that an apple is falling simply because it appears near a tree, relying on prior expectations rather than visual evidence.
These challenges highlight the need for going beyond raw video pixels and high-level textual supervision. Instead of training VLMs from scratch, which is infeasible and very costly in time and computational resource, one more data-efficient way is to augment VLMs with structured spatial and temporal representations that capture fine-grained object motions, interactions, and geometry. Recent advances in motion grounding [16], visual grounding [34, 37], detailed temporal modeling [64], and action analysis [17, 23] demonstrate that smaller specialized models can provide highly accurate spatial–temporal cues. However, existing approaches rarely explore how such modules can be systematically integrated into VLMs to elevate higher-level physics reasoning. Meanwhile, prior physics-oriented datasets [36, 29, 5] primarily rely on coarse annotations, limiting their capacity to induce deep, mechanism-level understanding.
@@ 2 Related Work
Physical Reasoning and Abnormality Detection in Videos: Despite rapid progress in AIGC video generation, generated videos still often contain prompt mismatches and violations of physics or commonsense, from earlier models such as LaVIE [54], SORA [7], and CogVideoX [63] to more recent systems including VEO3 [51], Wan2.2 [52], and COSMOS [1]. This has motivated growing work on abnormality evaluation and mitigation. VideoScore [22] trains synthetic evaluators aligned with human judgment, while recent benchmarks [29, 45, 44, 15, 36, 66] support systematic evaluation of physics and commonsense reasoning. WorldScore [18] provides a unified protocol over controllability, quality, and dynamics, and CRAVE [50] studies content-rich assessment through text–temporal fusion and motion-fidelity modeling. On the mitigation side, prior work has explored physics-based cues [24], chain-of-thought reasoning in PhyT2V [59], action-centric evaluation and correction in VideoPhy-2 [5], and motion-aware plausibility judgment in TRAVL [43]. Still, reliably identifying and addressing physical and commonsense failures remains an open challenge.
Spatial, Temporal, and Motion Understanding in VLMs: Recent work has improved video understanding in VLMs along spatial, temporal, and motion dimensions. For spatial reasoning, SpatialRGPT [12], VG-LLM [68], and LayoutVLM [49] strengthen 3D perception through depth cues, video-grounded geometry, and layout optimization. For temporal modeling, SlowFast-LLaVA [58] captures both local detail and long-range dynamics, TVS [19] improves alignment with temporally relevant segments, and T* [64] studies long-video retrieval via spatial search. Other efforts focus on motion and action understanding, including OpenMixer [6], MotionSight [17], and GroundMoRe [16]. However, physics-centric reasoning remains underexplored, as it requires jointly integrating spatial, temporal, and motion cues with physical commonsense.
Video Understanding in VLMs: Recent VLMs, including Qwen-VL [3] and InternVL [70], have improved general video understanding, yet remain limited on complex spatial, temporal, and long-video reasoning [35]. LongVLM [56] improves long-video modeling through hierarchical token merging, while VideoMind [42] uses a role-based agent with Chain-of-LoRA for efficient temporal reasoning. Video-Holmes [13] highlights persistent failures on multi-segment reasoning, and EgoLife [61] supports long-context retrieval for egocentric video assistants. Other work addresses hallucination and reasoning, including MASH-VLM [2], MVoT [28], and large-scale 3D spatial VQA data in [10]. Reinforcement-based training has also shown promise, as in Video-R1 [20] and VideoChat-R1 [32].
@@ 3 MASS-Bench: A Motion-Grounded Physics Reasoning and Comprehension Benchmark
Figure 2: Data Exhibition of MASS-Bench. MASS-Bench provides two question types, including factual and critical-thinking, to evaluate physics-driven video understanding. For each video–question–answer pair, we supply rich motion-grounding annotations, including temporal segmentation, entity-level visual grounding, temporal profiles across the full video, and motion attributes such as first/last positions and 3D displacement vectors. These structured spatial–temporal cues transform complex physics-related perception into interpretable representations that support more reliable physical reasoning. Additional dataset details are provided in the supplementary material.
Data Collection: MASS-Bench is a free-form video question-answering dataset designed to strengthen VLMs’ physics understanding, built from real-world videos sourced from Motion-Sight [17] and ActivityNet [8], alongside AIGC videos collected from VideoPhy2 [5], VideoHallu [36], and generative models such as COSMOS [1], Wan2.2 [52], and Sora [7]. MASS-Bench consists of $4,350$ videos, composed by $8,361$ free-form video question-answering pairs related to physics-related comprehension tasks. The dataset includes both positive and negative demonstrations with detailed human annotations providing balanced coverage of labels and reasoning steps. This design enables VLMs to learn correct physical processes from real-world videos, while also recognizing artifacts and violations that commonly arise in AIGC videos but are rare in natural scenes.
@@
Categorization: To capture the spectrum of physics-related cognition required in MASS-Bench, we group all questions into five categories ordered from basic to advanced: Spatial Understanding (SU, 33.3%), Temporal Understanding (TU, 19.5%), Motion & Action Recognition (MAR, 14.4%), Physics Comprehension (PC, 15.6%), and Physical Abnormality Detection (PA, 17.1%). VLMs must first establish spatial and temporal awareness, then recognize motion patterns, before progressing to higher-level physics comprehension and real-world violation detection. The full category definitions and distribution are provided in Table 5 of the supplementary material.
Motion Grounding. Beyond the question, video, and ground-truth answer, each pair includes motion-grounding annotations: temporal segmentation indexed by frame ranges, entity-level visual grounding with bounding boxes, temporal profiles tracking each entity across the full video, and per-segment motion attributes (first/last 3D positions and motion vectors). These cues convert physics-intensive perceptual challenges into structured textual representations for more reliable reasoning. Full annotation details are provided in the supplementary material.
Figure 3: Overview of MASS: We use a model-agnostic framework to enhance video understanding with explicit spatial and motion awareness. Beyond standard visual transformer encoders that process video inputs (e.g., LLaVA-OneVision [26], Qwen2.5-VL [3]), we introduce a visual grounding module to better align queried entities with corresponding visual cues. Depth estimation captures spatial geometry, while motion tracking captures temporal dynamics across frames. These cues are fused into entity-centric motion traces and converted into motion-derived textual tokens that are appended to the textual input stream. During post-training, we keep the spatial–temporal encoders frozen, i.e., used off-the-shelf without parameter updates, and apply reinforcement fine-tuning (RFT) only to improve the LLM backbone’s comprehension of the additional multimodal information.
@@ 4 MASS: Model-Agnostic Approach
High-quality human annotations are essential for strengthening VLMs’ understanding of physics-related visual contexts, yet data alone cannot address their core limitations in modeling spatial layouts and motion dynamics. Effective physical reasoning requires isolating relevant cues from noisy video content. While one solution is to retrain models with heavy preprocessing pipelines, this incurs substantial cost. Instead, we propose a more data-efficient alternative: augmenting VLMs with dedicated spatial and motion encoders that explicitly extract and represent key visual signals. As shown in Figure 3, our model-agnostic design combines lightweight architectural refinement with targeted training to enhance spatiotemporal perception, cross-modal alignment, and physics-aware reasoning, ultimately reducing hallucinations and improving physical comprehension.
@@ 4.1 Entity-Centric Visual Grounding
Understanding physics-related dynamics in videos often depends on accurately identifying and tracking the relevant entities across time. This remains challenging because videos contain dense spatial–temporal content, while language queries are often underspecified about which objects are relevant. Grounding is therefore critical for localizing queried entities, preserving their motion and spatial context, and filtering out irrelevant information. It also helps mitigate the information loss introduced by LLM-based VLMs such as LLaVA [40], which compress rich visual signals into a limited set of language-aligned tokens.
We begin by semantically extracting the key components from each probing question to narrow the target for visual grounding. Using Grounding-DINO [41], we detect bounding boxes for the queried entities and apply a dynamic temporal-resolution scheme that automatically selects an appropriate sub-sequence length based on the input video’s duration. This segmentation strategy increases temporal coherence within each chunk, improving object tracking and downstream processing while balancing accuracy and efficiency. It also preserves grounding consistency over time and reveals temporal artifacts such as sudden appearance or disappearance, which are common in AIGC videos. For each detected entity, we construct a time-aligned grounding profile. Finally, we apply SAM2 [46, 47] to generate segmentation masks for the grounded regions, enabling subsequent spatial–temporal analysis.